STOCHASTIC INTEGRAL AS A SPECTRAL INTEGRAL by Ljuban Dedić, Split, Croatia

Abstract. Dealing with stochastic integral with respect to L_{2} martingales with multidimensional time we prove that such an integral can be represented as a spectral integral with respect to a spectral measure generated by filtration. Mathematics subject classification: 28C20 Keywords: Wiener process, stochastic integral, spectral integral.